The electric flux density on a spherical surface r 5 b is the same for a point charge Q located at the origin and for charge Q uniformly distributed on surface r 5 a1a , b2. (a) Yes (c) Not necessarily (b) No


1
Expert's answer
2021-07-02T09:29:01-0400

Electric flux density="\smallint E.da=\smallint E.dacos\theta"

"\phi=\smallint E.dAcos\theta"

"\phi=E.A"

"\theta=0°"

r=b

Q charge distribution at sphere radius r=b

"\phi=\frac{kQ}{b^2}\times4\pi b^2"

"\phi=\frac{Q}{\epsilon_0}\rightarrow(1)"

r=a

Q charge distribution at sphere radius r=a

"\phi=\frac{kQ}{a^2}\times4\pi a^2"

"\phi=\frac{kQ}{a^2}\times4\pi a^2=\frac{Q}{\epsilon_0}\rightarrow(2)"

equation (1) and (2) are equal


Then electric flux density does not depend on size shape of sphere

Option (b) is correct option